Circular No. 6486

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1145–61 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association; and R. Remillard, Massachusetts Institute of Technology (MIT), report for the RXTE ASM Team at GSFC and MIT: “The Be/neutron binary star 4U 1145–61 (= 2S 1145–619) is experiencing an outburst. The outburst commenced on Sept. 29 UT rising suddenly to 43 ± 4 mCrab (2–12 keV), and daily detections continue through Oct. 8 with variable daily averages between 55 and 116 mCrab. This brightening is consistent with the 186.5-day period proposed by Warwick et al. (1985, Space Sci. Rev. 40, 429). However, during the previous predicted outburst at Mar. 31, any flare was significantly weaker than the current event. ASM

lightcurves are available via http://heasarc.gsfc.nasa.gov/docs/xte/
xte 1st.html.”

EXTREME-ULTRAVIOLET DETECTIONS OF COMETS J. P. Pye and R. G. West, Leicester University; and K. Dennerl, Max-

Planck-Institut für Extraterrestrische Physik, Garching, report the discovery of extreme-ultraviolet (EUV) emission from two comets in data acquired by the Wide-Field Camera (WFC) during the ROSAT all-sky survey. Accompanying x-ray emission has previously been reported by Dennerl et al. (IAUC 6413). Significant EUV detections have been made in the following observing intervals: C/1990 K1 (Levy), 1990 Sept. 6.50–7.43 UT, 1991 Jan. 9.96–11.96, and Jan. 17.36–18.02; C/1990 N1 (Tsuchiya-Kiuchi), 1990 Nov. 18.16–20.70. All detections were made in the WFC S1 filter, nominal passband 90–206 eV (14–6 nm). The upper-limit countrates derived for the WFC S2 filter (passband 62–110 eV; 20–11 nm), when compared with the S1 measurements, yield spectral parameters consistent with those previously reported for some comets (IAUC 6394, 6404).

COMET C/1996 E1 (NEAT) Total (visual) magnitude estimates (L = reflector): Aug. 4.91 UT,

11.3 (R. J. Bouma, Groningen, The Netherlands, 0.25-m L); 23.46, 11.7 (A. Hale, Cloudcroft, NM, 0.41-m L); Sept. 7.91, 10.9 (A. Baransky, Kiev, Ukraine, 0.11-m L); 17.13, 12.2 (G. W. Kronk, Troy, IL, 33.3-cm L); Oct. 3.91, 14.5 (H. Mikuž, Ljubljana, Slovenia, 0.36-m L + CCD + V filter).

1996 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green